# Containerization in Multi-Cloud Environment: Roles, Strategies, Challenges, and Solutions for Effective Implementation

### v1

@@ 3.3. Pattern and Strategies (RQ2)
Categories Subcategories
Patterns
ID
Cloud Architectural Patterns and Models Architecture Patterns
Microservice Architecture
S02, S09, S16, S77, S80, S24, S31, S33, S68
Service-Oriented Architecture (SOA)
S51, S52, S86
Multitier Architecture, Layered Architecture
S02, S04, S54
Client-Server Architecture
S03, S54
Component-based Architecture
S22
Event-driven Architecture
S44
Multi-agent Architecture
S74, S48
Master-worker Nodes Architecture
S33
Netflix Zuul-based Seeker Component
S17
Multi-tenant Cloud Service Architecture
S17
Stateful Engine Architecture
S17
Communication and Networking Patterns
Service Mesh
S05
Service Chaining (SC)
S50
Multi-Cloud APIs
S71
Third-Party APIs
S71
Service Request Broker
S07
Publish and Subscribe Communication Protocol
S45
Synchronous Network Communication Protocols
S01
Asynchronous Network Communication Protocols
S01
Network Function Virtualization (NFV)
S50
Service Discovery
S76
Deployment Pattern
Blue-Green deployment
S27, S30
Object store service
S76
Multi-cloud Deployment Model
S10
Distributed Deployment Model
S10
Accelerated Particle Swarm Optimization for Deployment
S16
Service Models
Platform-as-a-Service (PaaS)
S24, S41
Software as a Service
S79
Infrastructure-as-a-Service (IaaS)
S24
Function-as-a-Service (FaaS)
S41
Cloud Management and Resource Allocation Strategies Multi-Cloud Management Strategies
Multi-cloud computing strategy
S78, S81
Hybrid cloud architecture
S78
Hybrid/Multi-cloud Approach
S05, S83
Introducing multi-cloud Middleware
S52, S53
MAPE-K control loop
S46, S48
Connecting to multiple cloud service providers
S76
Multi-cloud load balancing
S23, S76
Multi-Cloud BP provisioning
S84
Cross-level orchestration of cloud services
S84
Cross-level monitoring and adaptation of BPs
S84
Container Management Strategies
Linux Container (LXC) project
S24
Container engine
S24
Docker Container Images
S24
Lightweight virtualization
S24
Portable application packaging
S24
One-container-per-app approach
S24
Data volumes and data volume containers
S24
Edge Computing and IoT Strategies
Mobile Edge Computing (MEC)
S22, S70, S43, S54
Fog computing
S32
Edge Services
S35
Connecting IoT Edge Devices to the Cluster
S05
Cloud of Things (CoT)
S05
Resource Management Strategies
Resource allocation based on different SLA levels
S23
Container Orchestration, Kubernetes, Minikube
S24, S29, S33
Cloud Security and Resilience Strategies Security and Resiliency Patterns
Security-by-Design approach
S51
Encryption of files
S71
Standardized APIs for file transfer
S71
Centralized/External ACM Service
S71
Distributed ACM Service
S71
Dynamic Switching between Authentication Methods
S71
ACM based on Signed URLs
S71
Attribute-Based Encryption and Signature
S75
Secure data sharing
S75
Data access control
S75
Local encryption and signing
S75
Authorized Tokens
S75
Byzantine quorum protocol
S75
Fault-tolerance Strategies
Fault-tolerance with Redundant Engines
S17
Multi-cloud systems fault-tolerant workflow
S67
RAFT Consensus Algorithm
S13
Cloud Migration Strategies
Service-oriented Migration
S15
Application-centric Migration
S15
Image-based Migration
S15
Migration to a Virtualized Container
S15
@@ 3.4. Quality Attributes and Tactics (RQ3)
Table 10 provides a comprehensive overview of the quality attributes and related tactics for containerized applications in multi-cloud environment. It is based on data extracted from 86 selected studies, which are organized following the ISO 25010 standard. This standard defines high-level QAs along with their sub-characteristics. Table 10 lists nine QAs, namely Performance (Efficiency), Security, Compatibility, Scalability, Reliability, Portability, Flexibility, Maintainability, and Usability. Each QA has its related terms, associated study numbers, and corresponding tactics to achieve them.
In Table 10, five tactics are provided for each quality attribute, due to the limited space. The complete set of 195 identified tactics is available in the provided dataset (Waseem et al., 2024), presenting a comprehensive map of tactics with QAs. However, it is important to note that these tactics are scattered across different studies, and only a few studies explicitly mentioned these tactics for specific QA.
Below we briefly discuss how each QA can be achieved by applying the mentioned tactics.
@@
Performance (Efficiency): The quality attribute of performance (or efficiency) in containerized applications can be achieved by employing tactics such as efficient resource optimization and allocation, use of ML techniques for performance optimization, utilization of container technology, location-aware service brokering, and performance-oriented Service Level Agreements (SLA). For instance, optimizing resources ensures that the system uses the minimum possible resources while delivering the required output. Using ML techniques for performance optimization can enhance system performance by learning and adjusting the operational parameters. Similarly, utilizing container technology can help maintain optimal performance by encapsulating the application and its environment.
@@
Security: Security can be enhanced using tactics like encryption and strong authentication mechanisms, implementation of firewalls, VPNs, or SDNs for network security, utilization of ML techniques to detect and prevent security threats and attacks, deployment of applications on diverse cloud providers, and consideration of users’ security specifications. For example, encryption provides a secure way of transmitting data, and ML can detect unusual behavior that could signal a threat or attack.
@@
Compatibility: The tactics that can be employed to achieve compatibility include implementation of lightweight communication protocols and modes, standardization of interfaces for seamless integration, componentization and modular design for interoperability, prototyping and exploration of interoperability approaches for multi-cloud deployment, and interoperability standardization and federation between clouds. Standardizing interfaces can ensure that different software components can interact with each other seamlessly, enhancing compatibility.
@@
Scalability: Scalability can be achieved through tactics like combining containerization and microservices for enhanced scalability, implementation of elastic resource allocation, enabling migration between multi-cloud services, combining container-based deployment with runtime monitoring and optimization, and leveraging orchestration, federated networks, and geographic placement. Elastic resource allocation ensures that the system can seamlessly scale up or down in response to changing demands.
@@
Reliability: Tactics such as deploying redundant engines for fault tolerance, distributing resources and replicating applications for improved response time, enforcing redundancy and distributed services for availability, deploying parallel search and multi-cloud distribution, and building fault-tolerant systems can enhance system reliability.
@@
Portability: Portability can be achieved using tactics such as virtual machine-based packaging, Docker image-based packaging, cloud-portable containerization, dynamic cross-level adaptation and provisioning, and adaptive rule-based system modification. These tactics ensure that an application can be easily transferred from one computing environment to another.
@@
Flexibility: Flexibility can be improved through tactics like service-oriented architecture, on-demand dynamic allocation of resources across different cloud platforms, containers, container orchestration, model-driven development, risk analysis, and cross-cloud service orchestration. Using multi-cloud environment, for instance, offers the flexibility to choose services from different providers as per specific needs.
@@
Maintainability: Tactics like microservices for easier maintenance, Infrastructure Provisioning as Code (IaC), containerization and image management, configuration management and templating, version control and change management, and continuous integration and deployment automation can enhance maintainability. For example, IaC allows developers to manage infrastructure more efficiently and minimize human errors, thus increasing maintainability.
@@
Usability: Usability can be enhanced by employing tactics like a feedback-loop controller for multi-cloud infrastructure, container orchestration (e.g., Kubernetes), user-centric interface design, responsive and adaptive user experience, accessible design and compliance, and error handling and feedback mechanisms. For instance, a feedback-loop controller is effective in a multi-cloud context as it dynamically adjusts resources across different cloud platforms based on user interaction, ensuring a seamless and responsive user experience tailored to the multi-cloud setup.
@@
Performance (Efficiency) Time behavior, Resource utilization, Capacity, Throughput, Response time S01, S02, S03, S06, S09, S14, S15, S16, S20, S21, S23, S25, S27, S29, S34, S35, S36, S37, S38, S39, S41, S43, S44, S47, S49, S52, S54, S57, S58, S59, S60, S62, S64, S65, S66 Efficient resource optimization and allocation
@@
Security Confidentiality, Authenticity, Access control, Authorization, Privacy, Trustworthiness, Integrity S01,S05, S06, S09, S15, S23, S24, S25, S29, S34, S37, S38, S40, S44, S47, S52, S54, S55, S57, S59, S61, S64, S65, S66, S69, S71, S72, S73, S77, S86 Encryption and strong authentication mechanisms
@@
Compatibility Interoperability, Co-existence S15, S16, S21, S24, S27, S42,S44, S45, S49, S63, S64, S71, S77, S80, S82 Implementing lightweight communication protocols and modes
@@
Scalability Capacity, Extensibility, Elasticity, Throughput, Responsiveness S01, S08, S09, S14, S15, S16, S17, S18,S19, S21, S22, S24, S25, S27, S29, S31, S35, S41, S42, S44, S45, S54, S57, S62, S66, S69, S72, S74, S76, S77, S82, S84
@@
Reliability Maturity, Availability, Fault tolerance, Recoverability S01, S02, S04, S10, S13, S17,S19, S20, S22, S25, S27, S32, S33, S35, S37, S44, S47, S51,S53, S54, S55, S56, S57, S60, S64, S65, S66, S67, S71, S73, S74, S75, S82 Deploying Redundant Engines for Fault-Tolerance
@@
Portability Installability, Replaceability, Adaptability S18, S24, S25, S31, S62, S77 Virtual Machine-Based Packaging
@@
Flexibility Functional completeness, Functional correctness, Functional appropriateness S16, S21, S27, S31, S41, S61, S60, S70, S76, S84 Service-oriented Architecture
@@
Maintainability Modularity, Reusability, Analyzability, Modfiabality, Testability S08, S17, S25, S27, S62, S74, S82 Microservices, easier maintenance
@@
Usability Appropriateness recognizability, Learnability, Operability , User interface aesthetics, Accessibility S09, S17, S27, S27, S37, S74 Feedback-loop controller for multi-cloud Infrastructure

### v2

@@ 3.3. Pattern and Strategies (RQ2)
Categories Subcategories Patterns ID
Cloud Architectural Patterns and Models Architecture Patterns Microservice Architecture (Jha et al., 2019; Verdugo et al., 2017; Ouyang et al., 2022; Ranjbar et al., 2017; Quint and Kratzke, 2018; Pahl, 2015; Mittal and Risco-Martín, 2017; Sami and Mourad, 2020; Guerrero et al., 2018a; Centofanti et al., 2024; Hossain et al., 2023; Tiwari and Sharma, 2023)
Service-Oriented Architecture (SOA) (Buyya and Barreto, 2015; Casola et al., 2017; Gundu et al., 2020; Merlino et al., 2024)
Multitier Architecture, Layered Architecture (Jha et al., 2019; Ramos et al., 2021; Bucur et al., 2018; Barati et al., 2023; Horchulhack et al., 2024; Liu and Qin, 2024; Tundo et al., 2024)
Client-Server Architecture (Hua et al., 2020; Bucur et al., 2018)
Component-based Architecture (Wong et al., 2019)
Event-driven Architecture (Wurster et al., 2018)
Multi-agent Architecture (Mfula and Nurminen, 2018; Saif et al., 2023)
Master-worker Nodes Architecture (Sami and Mourad, 2020)
Netflix Zuul-based Seeker Component (Slominski et al., 2015)
Multi-tenant Cloud Service Architecture (Slominski et al., 2015)
Stateful Engine Architecture (Slominski et al., 2015)
Distributed Multi-Cloud Native Architecture (Alonso et al., 2023a)
Plugin-Based Deployment Automation Architecture (Harzenetter et al., 2023)
Cloud-Edge Integrated Robotics Architecture (Huang and Wang, 2023)
Hybrid Malware Detection Architecture (Wang et al., 2023)
Cloudfront-Enhanced Container Architecture (Kataru et al., 2023)
AI-Driven Edge-Cloud Architecture (Benmerar et al., 2023)
Metadata-Driven Architecture (Orzechowski et al., 2023)
Blockchain-Based Architecture (Ahmad and Aujla, 2023)
Cross-Layered Edge-Cloud Architecture (Sofia et al., 2023; Santos et al., 2023)
Communication and Networking Patterns Service Mesh (Gattobigio et al., 2022)
Service Chaining (SC) (Bhamare et al., 2015)
Multi-Cloud APIs (Graupner et al., 2015a)
Third-Party APIs (Graupner et al., 2015a)
Service Request Broker (Zhang et al., 2021)
Publish and Subscribe Communication Protocol (Trihinas et al., 2015)
Synchronous Network Communication Protocols (Almeida et al., 2014)
Asynchronous Network Communication Protocols (Almeida et al., 2014)
Network Function Virtualization (NFV) (Bhamare et al., 2015)
Service Discovery (Pop et al., 2016)
Blockchain-Based Identity and Access Management (IAM) Architecture (Fugkeaw, 2023)
Federated Cloud-Edge Architecture (Roy and Kalafatis, 2023)
Deployment Pattern black-Green deployment (Alonso et al., 2019; Sheridan et al., 2017)
Object store service (Pop et al., 2016)
Multi-cloud Deployment Model (Sabbioni et al., 2020)
Distributed Deployment Model (Sabbioni et al., 2020)
Multi-Cloud Deployment Architecture (Petrovski and Gusev, 2024; Pandi et al., 2023)
Decentralized Orchestration Architecture (Sofia et al., 2024)
Policy-Driven Orchestration Architecture (Lee and Nam, 2023)
DevOps-Oriented Architecture (Solayman and Qasha, 2023)
Service Models Platform-as-a-Service (PaaS) (Pahl, 2015; Shi et al., 2020)
Software-as-a-Service (Serhiienko and Spillner, 2018)
Infrastructure-as-a-Service (IaaS) (Pahl, 2015)
Function-as-a-Service (FaaS) (Pahl, 2015; Ramesh et al., 2023)
Multi-Cloud Orchestration Architecture (Pandi et al., 2023)
Serverless Container-Oriented Architecture (Shi et al., 2020)
Cloud Management and Resource Allocation Strategies Multi-Cloud Management Strategies Multi-cloud computing strategy S78, S81
Hybrid cloud architecture (Wang et al., 2019)
Hybrid/Multi-cloud Approach (Gattobigio et al., 2022; Wahab et al., 2016)
Introducing multi-cloud Middleware (Casola et al., 2017; Acquaviva et al., 2017)
MAPE-K control loop (Fadda et al., 2019; Saif et al., 2023)
Connecting to multiple cloud service providers (Pop et al., 2016)
Multi-cloud load balancing (Wang et al., 2020; Pop et al., 2016)
Multi-Cloud BP provisioning (Kritikos et al., 2019)
Cross-level orchestration of cloud services (Kritikos et al., 2019)
Cross-level monitoring and adaptation of BPs (Kritikos et al., 2019)
Container Management Strategies Linux Container (LXC) project (Pahl, 2015)
Container engine (Pahl, 2015)
Docker Container Images (Pahl, 2015)
Lightweight virtualization (Pahl, 2015)
Portable application packaging (Pahl, 2015)
One-container-per-app approach (Pahl, 2015)
Data volumes and data volume containers (Pahl, 2015)
Edge Computing and IoT Strategies Mobile Edge Computing (MEC) (Wong et al., 2019; Lazuka et al., 2022; Dreibholz et al., 2019; Bucur et al., 2018)
Fog computing (Al-shammari and Alwan, 2018)
Edge Services (Carpio et al., 2020)
Connecting IoT Edge Devices to the Cluster (Gattobigio et al., 2022)
Cloud of Things (CoT) (Gattobigio et al., 2022)
Multi-Access Edge Computing (MEC) Architecture (Santos et al., 2023)
Cross-Layered Edge-Cloud Architecture (Sofia et al., 2023)
Cloud-Edge Integrated Robotics Architecture (Huang and Wang, 2023)
AI-Driven Edge-Cloud Architecture (Benmerar et al., 2023)
Resource Management Strategies Resource allocation based on different SLA levels (Wang et al., 2020)
Container Orchestration, Kubernetes, Minikube (Pahl, 2015; Gupta et al., 2021; Sami and Mourad, 2020)
Rate-Based Stream Processing Architecture (Nguyen and Chien, 2023)
Multi-Agent Resource Optimization Architecture (Nagarajan et al., 2023)
Cloud Security and Resilience Strategies Security and Resiliency Patterns Security-by-Design approach (Buyya and Barreto, 2015)
Encryption of files (Graupner et al., 2015a)
Standardized APIs for file transfer (Graupner et al., 2015a)
Centralized/External ACM Service (Graupner et al., 2015a)
Distributed ACM Service (Graupner et al., 2015a)
Dynamic Switching between Authentication Methods (Graupner et al., 2015a)
ACM based on Signed URLs (Graupner et al., 2015a)
Attribute-Based Encryption and Signature (Shen et al., 2017)
Secure data sharing (Shen et al., 2017)
Data access control (Shen et al., 2017)
Local encryption and signing (Shen et al., 2017)
Authorized Tokens (Shen et al., 2017)
Byzantine quorum protocol (Shen et al., 2017)
Blockchain-Driven SLA Management Architecture (Neeraj et al., 2023)
Hybrid Malware Detection Architecture (Wang et al., 2023)
Fault-tolerance Strategies Fault-tolerance with Redundant Engines (Slominski et al., 2015)
Multi-cloud systems fault-tolerant workflow (Tang, 2021)
RAFT Consensus Algorithm (Naik, 2016a)
Cloud Migration Strategies Service-oriented Migration (Pham and Pham, 2015)
Application-centric Migration (Pham and Pham, 2015)
Image-based Migration (Pham and Pham, 2015)
Migration to a Virtualized Container (Pham and Pham, 2015)
@@ 3.4. Quality Attributes and Tactics (RQ3)
Table 10 provides a comprehensive overview of the quality attributes and related tactics for containerized applications in multi-cloud environment. It is based on data extracted from 121 selected studies, which are organized following the ISO 25010 standard. This standard defines high-level QAs along with their sub-characteristics. Table 10 lists nine QAs, namely Performance (Efficiency), Security, Compatibility, Scalability, Reliability, Portability, Flexibility, Maintainability, and Usability. Each QA has its related terms, associated study numbers, and corresponding tactics to achieve them. In Table 10, due to limited space, we provide only five tactics for each QA. The complete set of 70 tactics is listed in the provided dataset (Waseem et al., 2024). However, it is important to note that these tactics are scattered across different studies. Below we briefly discuss how each QA can be achieved by applying the mentioned tactics.
@@
Performance (Efficiency): Performance or efficiency is the most dominant QA, frequently discussed in 60 (49.58%) of the selected studies, along with related terms or characteristics such as time behavior, resource utilization, and capacity. We also listed several tactics that can be used to achieve performance (or efficiency) in containerized applications, such as efficient resource optimization and allocation, use of machine learning techniques for performance optimization, utilization of container technology, location-aware service brokering, and performance-oriented Service Level Agreements (SLA). For instance, optimizing resources ensures that the system uses the minimum possible resources while delivering the required output. Using ML techniques for performance optimization can enhance system performance by learning and adjusting the operational parameters. Similarly, utilizing container technology can help maintain optimal performance by encapsulating the application and its environment. These results also indicate that performance is a critical QA, with a strong focus in containerized applications, and the majority of studies advocate for advanced tactics to achieve optimal performance.
@@
Security: This QA has been discussed in 38 studies along with various tactics. We also identified several characteristics of security that have been highlighted in the selected studies, such as confidentiality, authenticity, access control, authorization, privacy, trustworthiness, and integrity. Some of the identified tactics to improve the security of containerized applications from the selected studies include encryption and strong authentication mechanisms, implementation of firewalls, VPNs, or SDNs for network security, utilization of ML techniques to detect and prevent security threats and attacks, deployment of applications on diverse cloud providers, and consideration of users’ security specifications. For example, encryption provides a secure way of transmitting data, and ML can detect unusual behavior that could signal a threat or attack.
@@
Compatibility: This QA has been reported in 15 studies, along with characteristics such as interoperability and co-existence. Several tactics that can be employed to achieve compatibility include the implementation of lightweight communication protocols and modes, standardization of interfaces for seamless integration, componentization and modular design for interoperability, prototyping and exploring interoperability approaches for multi-cloud deployment, and interoperability standardization and federation between clouds. Standardizing interfaces can ensure that different software components can interact with each other seamlessly, enhancing compatibility. The results suggest that compatibility is essential for ensuring smooth integration and operation in multi-cloud environment.
@@
Scalability: Scalability refers to the ability to expand capacity or performance without losing the efficiency or functionality of a software system. We identified 40 studies that report scalability along with other characteristics like capacity, extensibility, elasticity, and throughput. We also identified ten tactics, five of which are listed in Table 10, to achieve scalability. For example, scalability can be achieved through tactics like combining containerization and microservices for enhanced scalability, implementing elastic resource allocation, enabling migration between multi-cloud services, combining container-based deployment with runtime monitoring and optimization, and leveraging orchestration, federated networks, and geographic placement. Elastic resource allocation ensures that the system can seamlessly scale up or down in response to changing demands.
@@
Reliability: This QA is about the system’s correct operation and consistent performance without failure over a specified period, while ensuring the availability of the system. We identified XX studies that report on reliability, along with characteristics such as maturity, availability, and fault tolerance. We identified nine tactics overall (see Table 10 and the QA and Tactics Sheet in (Waseem et al., 2024)) to achieve reliability from the selected studies. These include deploying redundant engines for fault tolerance, distributing resources and replicating applications for improved response time, enforcing redundancy and distributed services for availability, deploying parallel search and multi-cloud distribution, and building fault-tolerant systems to enhance system reliability.
@@
Portability: Portability refers to a system’s ability to function correctly across different platforms (e.g., operating systems, hardware configurations). We identified 6 studies that report on portability, along with characteristics such as installability, replaceability, and adaptability, as well as 5 tactics that can help achieve portability. For example, portability can be achieved using tactics such as virtual machine-based packaging, Docker image-based packaging, cloud-portable containerization, dynamic cross-level adaptation and provisioning, and adaptive rule-based system modification. These tactics ensure that an application can be easily transferred from one computing environment to another.
@@
Flexibility: This QA represents the system’s capacity to accommodate new features based on user requirements with minimal effort. We identified 10 studies (see Table 10 and the QA and Tactics Sheet in (Waseem et al., 2024)) that report this QA, along with five tactics that can help achieve it. For example, flexibility can be improved through tactics such as service-oriented architecture, on-demand dynamic allocation of resources across different cloud platforms, containers, container orchestration, model-driven development, risk analysis, and cross-cloud service orchestration. Using multi-cloud environment, for instance, offers the flexibility to choose services from different providers as per specific needs.
@@
Maintainability: This QA has been reported in 7 studies, along with several characteristics such as modularity, reusability, and analyzability. We also identified 6 tactics from these studies that can help achieve maintainability. For example, tactics like microservices for easier maintenance, Infrastructure Provisioning as Code (IaC), containerization and image management, configuration management and templating, version control and change management, and continuous integration and deployment automation can enhance maintainability. For example, IaC allows developers to manage infrastructure more efficiently and minimize human errors, thus increasing maintainability.
@@
Usability: This QA refers to the ease with which users can interact with a system to perform various operations. We identified 6 studies that report on usability, along with several characteristics such as learnability, and operability, as well as 6 other tactics. According to the selected studies, usability can be enhanced by employing tactics such as a feedback-loop controller for multi-cloud infrastructure, container orchestration (e.g., Kubernetes), user-centric interface design, responsive and adaptive user experience, accessible design and compliance, and error handling and feedback mechanisms. For instance, a feedback-loop controller is effective in a multi-cloud context as it dynamically adjusts resources across different cloud platforms based on user interactions, ensuring a seamless and responsive user experience tailored to the multi-cloud setup.
@@
Performance (Efficiency) Time behavior, Resource utilization, Capacity, Throughput, Response time (Almeida et al., 2014; Jha et al., 2019; Hua et al., 2020; Deng et al., 2022; Verdugo et al., 2017; Chen et al., 2022; Pham and Pham, 2015), (Ouyang et al., 2022; Ferry et al., 2014; Cui et al., 2016; Wang et al., 2020; Gao et al., 2017; Alonso et al., 2019; Gupta et al., 2021), (Rios et al., 2017; Carpio et al., 2020; Lorenzo et al., 2019; Dimitrijevic et al., 2019; Pandey et al., 2022; Altran et al., 2022; Shi et al., 2020), (Dreibholz et al., 2019; Wurster et al., 2018; Libardi et al., 2014; Harwalkar et al., 2019b; Casola et al., 2017; Bucur et al., 2018; Tricomi et al., 2017), (Sitaram et al., 2018; Rodigari et al., 2021; Janarthanan et al., 2018; Rahman and Lama, 2019; Harwalkar et al., 2019a; Zeck and Bouroudjian, 2017; Aral et al., 2020) Efficient resource optimization and allocation
@@
Security Confidentiality, Authenticity, Access control, Authorization, Privacy, Trustworthiness, Integrity (Almeida et al., 2014; Gattobigio et al., 2022; Deng et al., 2022; Verdugo et al., 2017; Pham and Pham, 2015; Wang et al., 2020; Pahl, 2015), (Gao et al., 2017; Gupta et al., 2021; Rios et al., 2017; Dimitrijevic et al., 2019; Pandey et al., 2022; Bélair et al., 2019; Wurster et al., 2018), (Libardi et al., 2014; Casola et al., 2017; Bucur et al., 2018; Bolivar et al., 2018; Tricomi et al., 2017; Rodigari et al., 2021; Di Modica et al., 2019), (Harwalkar et al., 2019a; Zeck and Bouroudjian, 2017; Aral et al., 2020; Kiss, 2018; Graupner et al., 2015a; Alaluna et al., 2017; Guechi and Maamri, [n. d.]; Ranjbar et al., 2017; Gundu et al., 2020) Encryption and strong authentication mechanisms
@@
Compatibility Interoperability, Co-existence (Pham and Pham, 2015; Ouyang et al., 2022; Cui et al., 2016; Pahl, 2015; Alonso et al., 2019; Baresi et al., 2016; Wurster et al., 2018), (Trihinas et al., 2015; Harwalkar et al., 2019b; Pellegrini et al., 2017; Harwalkar et al., 2019a; Graupner et al., 2015a; Ranjbar et al., 2017; Quint and Kratzke, 2018; Ferry et al., 2013) Implementing lightweight communication protocols and modes
@@
Scalability Capacity, Extensibility, Elasticity, Throughput, Responsiveness (Almeida et al., 2014; Pham et al., 2015; Verdugo et al., 2017; Chen et al., 2022; Pham and Pham, 2015; Ouyang et al., 2022; Slominski et al., 2015), (Pai and Kumar, 2021; Alhamazani et al., 2014; Cui et al., 2016; Wong et al., 2019; Pahl, 2015; Gao et al., 2017; Alonso et al., 2019), (Gupta et al., 2021; Mittal and Risco-Martín, 2017; Carpio et al., 2020; Shi et al., 2020; Baresi et al., 2016; Wurster et al., 2018; Trihinas et al., 2015), (Bucur et al., 2018; Tricomi et al., 2017; Rahman and Lama, 2019; Aral et al., 2020; Kiss, 2018; Alaluna et al., 2017; Mfula and Nurminen, 2018), (Pop et al., 2016; Ranjbar et al., 2017; Ferry et al., 2013; Kritikos et al., 2019)
@@
Reliability Maturity, Availability, Fault tolerance, Recoverability (Almeida et al., 2014; Jha et al., 2019; Ramos et al., 2021; Sabbioni et al., 2020; Naik, 2016a; Slominski et al., 2015; Alhamazani et al., 2014), (Ferry et al., 2014; Wong et al., 2019; Gao et al., 2017; Alonso et al., 2019; Al-shammari and Alwan, 2018; Sami and Mourad, 2020; Carpio et al., 2020), (Dimitrijevic et al., 2019; Wurster et al., 2018; Libardi et al., 2014; Buyya and Barreto, 2015; Acquaviva et al., 2017; Bucur et al., 2018; Bolivar et al., 2018), (del Castillo et al., 2013; Tricomi et al., 2017; Janarthanan et al., 2018; Harwalkar et al., 2019a; Zeck and Bouroudjian, 2017; Aral et al., 2020; Tang, 2021), (Graupner et al., 2015a; Guechi and Maamri, [n. d.]; Mfula and Nurminen, 2018; Shen et al., 2017; Ferry et al., 2013) Deploying Redundant Engines for Fault-Tolerance
@@
Portability Installability, Replaceability, Adaptability (Pai and Kumar, 2021; Pahl, 2015; Gao et al., 2017; Mittal and Risco-Martín, 2017; Rahman and Lama, 2019; Ranjbar et al., 2017) Virtual Machine-Based Packaging
@@
Flexibility Functional completeness, Functional correctness, Functional appropriateness (Ouyang et al., 2022; Cui et al., 2016; Alonso et al., 2019; Mittal and Risco-Martín, 2017; Shi et al., 2020; Di Modica et al., 2019; Janarthanan et al., 2018), (Lazuka et al., 2022; Pop et al., 2016; Kritikos et al., 2019) Service-oriented Architecture
@@
Maintainability Modularity, Reusability, Analyzability, Modfiabality, Testability (Pham et al., 2015; Slominski et al., 2015; Gao et al., 2017; Alonso et al., 2019; Rahman and Lama, 2019; Mfula and Nurminen, 2018; Ferry et al., 2013) Microservices, easier maintenance
@@
Usability Appropriateness recognizability, Learnability, Operability , User interface aesthetics, Accessibility (Verdugo et al., 2017; Slominski et al., 2015; Alonso et al., 2019, 2019; Dimitrijevic et al., 2019; Mfula and Nurminen, 2018) Feedback-loop controller for multi-cloud Infrastructure